OpenStudy (anonymous):

how do you evaluate the indefinite integral (arcsinx)^7/(√1-x^2) dx ? so tricky

OpenStudy (anonymous):

If it's\[\int\limits_{a}^{b}(\sin^{-1}(x)^{7}/(\sqrt{1-x^{2}}))dx\]Try u = arcsinx, du = (1/(sqrt(1-x^2))dx

OpenStudy (anonymous):

It's a simple power rule integral after that.

OpenStudy (anonymous):

would u mind giving me the answer i have to have an assignment done in exactly 4 minutes and its my 2nd last question!!

OpenStudy (anonymous):

It's just\[\int\limits_{a}^{b}u^{7}du\]where u = arcsin(x). Just solve it and substitute it back.